### v2

@@ Abstract
We have discovered the stellar counterpart to the ALFALFA Virgo 7 cloud complex, which has been thought to be optically dark and nearly star-free since its discovery in 2007. This 190 kpc long chain of enormous atomic gas clouds ( M) is embedded in the hot intracluster medium of the Virgo galaxy cluster but is isolated from any galaxy. Its faint, blue stellar counterpart, BC6, was identified in a visual search of archival optical and UV imaging. Follow-up observations with the Green Bank Telescope, Hobby-Eberly Telescope, and Hubble Space Telescope demonstrate that this faint counterpart is at the same velocity as the atomic gas, is actively forming stars, and is metal-rich (). We estimate its stellar mass to be only , making it one of the most gas-rich stellar systems known. Aside from its extraordinary gas content, the properties of BC6 are entirely consistent with those of a recently identified class of young, low-mass, isolated, and star-forming clouds in Virgo, that appear to have formed via extreme ram pressure stripping events. We expand the existing discussion of the origin of this structure and suggest NGC 4522 as a likely candidate, however, the current evidence is not fully consistent with any of our proposed progenitor galaxies. We anticipate that other “dark” gas clouds in Virgo may have similarly faint, star-forming counterparts. We aim to identify these through the help of an ongoing citizen science search of the entire cluster.
@@ 1 Introduction
Although some of the structures identified in the references above are still connected to their parent galaxies, some are detached, isolated, and apparently dark H i clouds with no clear point of origin. Such objects span several orders of magnitude in H i mass and are certainly not a single type of object. At the high-mass end, Kent et al. (2007) and Taylor et al. (2012) identified several dark H i clouds () in the direction of the Virgo cluster. At intermediate masses, the Five-hundred-meter Aperture Spherical Telescope and the Green Bank Telescope (GBT) have recently identified lone H i clouds that may be ancient objects (Zhou et al., 2023; Karunakaran & Spekkens, 2024, K. O’Neil private communication). At the low-mass end, Adams et al. (2013) identified a sample of optically dark, compact H i clouds, some of which could be very low mass halos in or near the Local Group.
When dealing with H i-only detections of local objects (e.g. km s) it is challenging to determine accurate distances, and these systems may not be as they first appear. Adams et al. (2015), Bellazzini et al. (2015), and Sand et al. (2015) all, almost simultaneously, realized that one of the Adams et al. (2013) objects (AGC 226067) did contain a small number of stars and was 10 farther away and 100 more massive than first thought. This object, now commonly referred to as SECCO 1, was found to be an isolated pocket of recent star formation (SF) embedded in the hot intra-cluster medium (ICM) of the Virgo cluster. Its rich metallicity (Beccari et al., 2017) and exclusively young stellar population (Sand et al., 2017) indicated that it must have formed recently from stripped gas, but there was no candidate parent galaxy within over 200 kpc.
@@
With the goal of better understanding the population statistics of these new objects we embarked on a uniform search for candidates covering the entire cluster (described in §2). When we began following up our initial candidates (§3), one (known as BC6) was found to be coincident (both in position and velocity) with the largest of the “dark” H i cloud complexes identified by Kent et al. (2007). Despite repeated study of this cloud complex (Kent et al., 2009; Sorgho et al., 2017; Minchin et al., 2019) no convincing stellar counterpart had been identified since its discovery. Here we show (§4) that it has a definite stellar counterpart that is over 20,000 times less massive than the H i, and is made up exclusively of stars that recently formed in situ. The stellar properties of this object closely match those of the recently identified blue blobs (Jones et al., 2022b). BC6 differs only in its exceptional H i content. In §5 we discuss the potential point of origin of this object, and its likely fate. We present our conclusions in §6.
@@ 3 Observations
BC6 lies in the direction of the Virgo cluster approximately 2.7 south of M 87 and 1.7 north of M 49. That is, roughly halfway been the Virgo A and B clouds (Binggeli et al., 1987). Typical cluster members in this direction have line of sight velocities in the range , but the scatter is several hundred km s beyond either side of this range (Mei et al., 2007). There is a bridge of diffuse X-ray emission that connects M 87 to M 49 (Böhringer et al., 1994), indicating that the hot ICM is higher density in this region than would be expected given its separation from the cluster core. BC6 lies roughly on this bridge.
@@ 3.2 Green Bank Telescope
Before it was realized that BC6 was the stellar counterpart to the H i structure identified by Kent et al. (2007), and that it therefore already had H i observations, it was included as a target in the GBT project. BC6 was observed in April 2023 with six 10 min on/off integrations with the VEGAS (Versatile GBT Astronomical Spectrometer) backend. The data were reduced using standard procedures in GBTIDL.555https://gbtidl.nrao.edu After smoothing to a spectral resolution of 5 km s, the resulting spectrum (Figure 2) has an rms noise of 1.77 mJy. H i line emission is strongly detected in the 9.1′ GBT beam along the BC6 line-of-sight between 475 km s – 625 km s.
@@ 3.3 ALFALFA
The Arecibo Legacy Fast ALFA (Arecibo L-band Feed Array), or ALFALFA, survey (Giovanelli et al., 2005; Haynes et al., 2011, 2018) mapped approximately 3000 sq deg of the sky for H i line emission, including the Virgo cluster. Kent et al. (2007) originally identified the H i cloud complex in a preliminary ALFALFA data set. These data have a channel width of 5 km s, a pixel size of 1′, and a spatial resolution of approximately 3.8′. We created our own version of the moment zero map of the H i emission in the cloud complex using the SoFiA (Serra et al., 2014, 2015) masking tool, adopting the standard smooth and clip algorithm with spatial smoothing kernels of 4.0 and 8.0′, with both no spectral smoothing and smoothing over two and four channels (10 and 20 km s). We set a detection threshold of 3.5 in order to include relatively faint emission, but coupled this with a high reliability threshold (95%) in order to ensure that only sources that are high S/N overall were retained. The moment zero map of the emission included in the SoFiA mask is shown in Figure 3 (left) as contours on a DECaLS optical image.
We estimate the H i mass of the entire structure by integrating all of the flux within the SoFiA mask and assuming a distance of 16.5 Mpc. This gives a value of . We note that this is about 0.1 dex higher than the value in Kent et al. (2009), however, in that work they considered each cloud component as a separate source, which may have led to some extended flux being missed. In addition, Minchin et al. (2019) used deeper Arecibo observations to estimate a total H i mass of , arguing that the additional flux (compared to ALFALFA) comes from low column density, extended emission.
@@ 3.4 Hubble Space Telescope imaging
HST Advanced Camera for Surveys (ACS) images of BC6666http://dx.doi.org/10.17909/h12e-gb42 (catalog DOI: 10.17909/h12e-gb42) were taken in February 2023 as part of the program GO-17267 (PI: M. Jones). The imaging used the Wide Field Channel of ACS with the F606W and F814W filters. Over the course of two orbits, four exposures were taken in each filter, resulting in total integration times of approximately 2000 s in F606W and F814W. A false color composite of these images is shown in Figure 3 (right), centered on BC6.
@@ 3.5 Hobby-Eberly Telescope optical spectroscopy
We observed the main component of BC6 (the largest ellipse in Figure 3, right) with the Blue spectrograph of the Low Resolution Spectrograph 2 (LRS2-B, Chonis et al., 2016) on the 10m HET (Ramsey et al., 1998; Hill et al., 2021) under clear dark skies on the night of April 19th, 2023 with a 3300 s exposure. The LRS2-B spectrograph has and an integral field unit which is 6″12″ with 0.6″ lenslet fibers on sky, giving an abundance of sky fibers around the relatively compact optical emission from BC6. We use the H line to determine which spaxels include emission from BC6 and extract all spaxels within 1.5″ of the peak of the source with the LRS2Multi extension to the Panacea data reduction pipeline.777https://github.com/grzeimann/LRS2Multi and https://github.com/grzeimann/Panacea Briefly, this data reduction includes the necessary detector characterizations, fiber extraction, wavelength calibration, mirror illumination correction, sky subtraction, and 1-D spectral source extraction including a treatment of differential atmospheric refraction in the sky and object spectra. The resulting spectrum is shown in Figure 5.
@@
The spectrum is fairly typical of an H ii region, with clear, albeit relatively weak, emission lines. Only the H line is high enough S/N to be fit with a Gaussian. Thus, this line alone was used to determine the redshift km s (with a barycentric correction applied). To obtain fluxes from the other lines we extract the integrated signal within a 2 region centered on the redshifted wavelength on each line. Here is the standard deviation of the Gaussian profile fit to the H line. Flux errors are estimated by summing the uncertainty values in the same wavelength range and dividing by the square root of the number of channels included.
@@
A reddening correction was calculated by assuming an intrinsic value of (for an H ii region at K; Osterbrock, 1989), an value of 3.1, and the Fitzpatrick (2004) extinction law. Based on the observed line ratio we calculated an extinction value of . This suggests that there is some internal extinction within BC6 as the Schlegel et al. (1998) dust map in this direction indicates minimal Galactic extinction (). Observed line fluxes were corrected for extinction (using the extinction model of Fitzpatrick, 2004) implemented in the Python package dust_extinction.888https://dust-extinction.readthedocs.io These fluxes and uncertainties are provided in Table 2. For consistency the quoted H flux does not correspond to that calculated with the Gaussian fit, but was instead calculated in the same manner as the other lines.
@@
With these line flux measurements we used two line ratio metrics to estimate the gas-phase oxygen abundance in BC6 (following Pettini & Pagel, 2004), N2=[Nii]/H and O3N2=([Oiii]/H)/([Nii]/H). We note that these ratios are designed to be almost independent of our extinction correction as [Oiii] and H occur at almost the same wavelength, as do [Nii] and H. We see (in Table 2) that both metrics give almost exactly the same oxygen abundance, despite each method having an intrinsic scatter of 0.2 dex.
@@ 4.2 Metallicity
Following the same rationale as Jones et al. (2022b), the above metallicity estimate implies that the progenitor galaxy of BC6 was likely in the stellar mass range , assuming it roughly follows the mass–metallicity relation (Andrews & Martini, 2013). The broad range is a reflection of the large uncertainty in the metallicity. Unfortunately, over this stellar mass range the typical gas fraction () of a field galaxy smoothly transitions over the range (e.g. Huang et al., 2012), thus the fact that BC6 is accompanied by 10 M of H i gas does little to help narrow down the progenitor’s stellar mass. However, it does mean that the progenitor must have lost a significant fraction, probably the majority, of its H i gas reservoir.
@@ 4.3 Stellar mass
To estimate the stellar mass of BC6 we follow the same approach as Jones et al. (2022b), that is, assume a constant SFR and continuously build up a synthetic stellar population (of matching metallicity) until the observed total F814W magnitude of BC6 is reached. Using this approach we estimate a stellar mass of M, which at the observed (NUV-based) SFR could have been built up over approximately 80 Myr. If we instead use and -band aperture photometry from the NGVS images and the scaling relations of Zibetti et al. (2009) and Taylor et al. (2011), we obtain estimates of and M, respectively.
@@
With a stellar mass of M, BC6 must have a gas fraction () of around 3,000 if we only consider the gas in the coincident cloud (C3; Figure 3, left), or over 20,000 if we include the whole H i cloud complex. This makes it one of, if not the most H i-rich extragalactic stellar systems known.
@@ 5.1 Point of origin
As mentioned by Minchin et al. (2019), tidal stripping does not seem to be a viable mechanism to form this structure. Not only is its isolation more difficult to explain with the low velocities (300 km s) involved in tidal stripping (e.g. Bournaud & Duc, 2006), but M represents a significant amount of H i gas for almost any galaxy and it is not possible to tidally remove a large fraction of a galaxy’s gas disk without also stripping stars. The exception to this would be if the progenitor was a particularly massive and H i-rich galaxy, however, that would also require a close passage from a similarly massive perturber to strip the gas, with neither remaining connected to the stripped gas in the present. Thus, the complete lack of stars over almost all of such a massive H i complex is a strong argument for ram pressure stripping and we will only consider the formation of BC6 via this mechanism.
@@
The H i cloud complex (Figure 3) is about 40′ across (190 kpc), making it plausible that the progenitor galaxy is a degree or more away. Unfortunately, despite there being no clear progenitor galaxy in the immediate vicinity of BC6, there are 10s of potential parent objects that are Virgo members within a few degrees. Both Kent et al. (2009) and Minchin et al. (2019) focused on NGC 4445 and NGC 4424 as these are only 30′ to the SW of BC6 and have similar radial velocities to the gas complex (346 km s and 437 km s, respectively; Albareti et al., 2017; Haynes et al., 2018). NGC 4424 is in the process of being ram pressure stripped (Chung et al., 2009; Sorgho et al., 2017), but its H i tail points to the SE, roughly parallel to the H i complex containing BC6, rather than towards it. In the case of NGC 4445 the –B-band diameter scaling relation of Jones et al. (2018) indicates that its expected original H i content (before falling into the Virgo cluster) would have been (and it still contains ; Haynes et al., 2018). The observed mass of the H i complex therefore represents almost its entire original H i content. Furthermore, in Sorgho et al. (2017), figure A1, NGC 4445 also appears to have a ram pressure tail (though much smaller than that of NGC 4424) which is pointing almost directly away from BC6 and the H i cloud complex. Thus, neither of these appear to be a viable candidate for the progenitor system.
@@
In simulations of high speed ram pressure stripping events (e.g. Kapferer et al., 2009; Tonnesen & Bryan, 2012, 2021; Göller et al., 2023), on scales much larger than a galaxy, the tails that are formed are usually relatively linear features. This means that for progenitor candidates that are within 1, but not aligned with the major axis of the H i complex, a larger than expected opening angle would be required to explain the geometry of the structure. It would perhaps still be possible to reproduce the observed geometry by involving clumpy structure within the ICM (the simulations assume the ICM is uniform and moving at a constant velocity). However, in an attempt to narrow our focus to the most likely candidates, we will now consider potential progenitors that are roughly NW and SE of BC6, such that the cloud geometry can readily be explained by a simple ram pressure stripping scenario. There are five potentially viable progenitors within a few degrees: NGCs 4380 and 4390 to the NW, and NGCs 4519, 4522, and 4535 to the SE. All five are massive enough to meet the mass–metallicity requirement from §4.2 and all currently contain some H i, detected in ALFALFA. The basic properties of these galaxies are summarized in Table 3.
Galaxy | /km s | /km s | Separation | Direction | [M/H] | ||
NGC 4380 | 963 | 463 | 64′ | NW | 8.51 | 9.40 | |
NGC 4390 | 1101 | 601 | 70′ | NW | 8.68 | 8.85 | -0.38 |
NGC 4424 | 437 | -63 | 40′ | W | 8.34 | 9.43 | 0.04 |
NGC 4445 | 353 | -147 | 27′ | W | 7.71 | 9.19 | |
NGC 4519 | 1216 | 716 | 87′ | SE | 9.47 | 9.33 | -0.18 |
NGC 4522 | 2329 | 1819 | 69′ | SE | 8.66 | 9.45 | -0.01 |
NGC 4535 | 1963 | 1463 | 116′ | SE | 9.57 | 9.93 | 0.06 |
@@
We see from the metallicity measurements that NGC 4519 and NGC 4522 are the closest matches to BC6 (which has ). However, there is no measurement for NGC 4380, and NGC 4535 could still be a match given the uncertainties. Thus, only NGC 4390 is strongly disfavored because of its metallicity. The predicted original H i mass of NGC 4390 also suggests that it never would have had enough H i to form the cloud complex associated with BC6. Furthermore, it appears NGC 4390 has not lost a significant amount of H i.999We note that the predicted original H i mass is usually intended as a statistical quantity and has considerable uncertainty for any individual galaxy. For example, the measured H i content of NGC 4519 exceeds its predicted original H i mass by over 0.1 dex. These values should be considered a rough guide only.
@@
NGC 4380 was mapped in the VLA Imaging of Virgo in Atomic gas project (VIVA; Chung et al., 2009), which found that it has a truncated H i disk. This indicates a past episode of ram pressure stripping, however, the direction of stripping is unknown.
@@
Unfortunately, VIVA did not map either NGC 4519 or NGC 4535. Both are detected in ALFALFA, but these data have an angular resolution of 4′. At this resolution NGC 4535 appears undisturbed, but the H i distribution of NGC 4519 is clearly extended to the NW (towards BC6). However, NGC 4519 has a neighbor immediately to the NW (NGC 4519A) separated by only 200 km s (Binggeli et al., 1985). Thus the disturbance in the H i distribution of NGC 4519 may be the result of either tides or ram pressure and with the resolution of the ALFALFA data there is little chance of distinguishing the two.

### v3

@@ Abstract
We have discovered the stellar counterpart to the ALFALFA Virgo 7 cloud complex, which has been thought to be optically dark and nearly star-free since its discovery in 2007. This 190 kpc long chain of enormous atomic gas clouds ( M⊙) is embedded in the hot intracluster medium of the Virgo galaxy cluster but is isolated from any galaxy. Its faint, blue stellar counterpart, BC6, was identified in a visual search of archival optical and UV imaging. Follow-up observations with the Green Bank Telescope, Hobby-Eberly Telescope, and Hubble Space Telescope demonstrate that this faint counterpart is at the same velocity as the atomic gas, is actively forming stars, and is metal-rich (). We estimate its stellar mass to be only , making it one of the most gas-rich stellar systems known. Aside from its extraordinary gas content, the properties of BC6 are entirely consistent with those of a recently identified class of young, low-mass, isolated, and star-forming clouds in Virgo, that appear to have formed via extreme ram pressure stripping events. We expand the existing discussion of the origin of this structure and suggest NGC 4522 as a likely candidate, however, the current evidence is not fully consistent with any of our proposed progenitor galaxies. We anticipate that other “dark” gas clouds in Virgo may have similarly faint, star-forming counterparts. We aim to identify these through the help of a citizen science search of the entire cluster.
@@ 1 Introduction
Although some of the structures identified in the references above are still connected to their parent galaxies, some are detached, isolated, and apparently dark H i clouds with no clear point of origin. Such objects span several orders of magnitude in H i mass and are certainly not a single type of object. At high-mass and intermediate masses Kent et al. (2007) and Taylor et al. (2012) identified several dark H i clouds () in the direction of the Virgo cluster and the Five-hundred-meter Aperture Spherical Telescope and the Green Bank Telescope (GBT) have recently identified lone H i clouds that may be ancient objects (Zhou et al., 2023; Karunakaran & Spekkens, 2024; O’Neil et al., 2024). At the low-mass end, Adams et al. (2013) identified a sample of optically dark, compact H i clouds, some of which could be very low mass halos in or near the Local Group.
When dealing with H i-only detections of local objects (e.g. km s-1) it is challenging to determine accurate distances, and these systems may not be as they first appear. Adams et al. (2015), Bellazzini et al. (2015), and Sand et al. (2015) all, almost simultaneously, realized that one of the Adams et al. (2013) objects (AGC 226067) did contain a small number of stars and was 10 farther away and 100 more massive than first thought. This object, now commonly referred to as SECCO 1, was found to be an isolated pocket of recent star formation (SF) embedded in the hot intra-cluster medium (ICM) of the Virgo cluster. Its rich metallicity (Beccari et al., 2017) and exclusively young stellar population (Sand et al., 2017) indicated that it must have formed recently from stripped gas, but there was no candidate parent galaxy within over 200 kpc.
@@
With the goal of better understanding the population statistics of these new objects we embarked on a uniform search for candidates covering the entire cluster (described in §2). When we began following up our initial candidates (§3), one (known as BC6) was found to be coincident (both in position and velocity) with the largest of the “dark” H i cloud complexes identified by Kent et al. (2007), ALFALFA Virgo 7. In a follow-up investigation, Kent et al. (2009) found that VCC 1357 overlaps with this H i structure, but was deemed to likely be in projection and not physically associated with the gas due to its color and dwarf spheroidal morphology. Another faint coincident object was noted by Kent et al. (2009), but more recent imaging clearly indicates that this is a distant background group of galaxies. Thus, no convincing stellar counterpart had been identified for this H i structure since its discovery. Here we show (§4) that ALFALFA Virgo 7 has a definite stellar counterpart that is over 20,000 times less massive than the H i, and is made up exclusively of stars that recently formed in situ. The stellar properties of this object closely match those of the recently identified blue blobs (Jones et al., 2022b). BC6 differs only in its exceptional H i content. In §5 we discuss the potential point of origin of this object, and its likely fate. We present our conclusions in §6.
@@ 3 Observations
BC6 lies in the direction of the Virgo cluster approximately 2.7∘ south of M 87 and 1.7∘ north of M 49. That is, roughly halfway been the Virgo A and B clouds (Binggeli et al., 1987). Typical cluster members in this direction have line of sight velocities in the range , but the scatter is several hundred km s-1 beyond either side of this range (Mei et al., 2007). There is a bridge of diffuse X-ray emission that connects M 87 to M 49 (Böhringer et al., 1994), indicating that the hot ICM is higher density in this region than would be expected given its separation from the cluster core. BC6 lies roughly on this bridge.
@@ 3.2 Green Bank Telescope
Before it was realized that BC6 was the stellar counterpart to the H i structure identified by Kent et al. (2007), and that it therefore already had H i observations, it was included as a target in the GBT project. BC6 was observed in April 2023 with six 10 min on/off integrations with the VEGAS (Versatile GBT Astronomical Spectrometer) backend. The data were reduced using standard procedures in GBTIDL.555https://gbtidl.nrao.edu After smoothing to a spectral resolution of 5 km s-1, the resulting spectrum (Figure 2) has an rms noise of 1.77 mJy. H i line emission is strongly detected in the 9.1′ GBT beam along the BC6 line-of-sight between 475 km s-1 – 625 km s-1.
@@ 3.3 ALFALFA
The Arecibo Legacy Fast ALFA (Arecibo L-band Feed Array), or ALFALFA, survey (Giovanelli et al., 2005; Haynes et al., 2011, 2018) mapped approximately 3000 sq deg of the sky for H i line emission, including the Virgo cluster. Kent et al. (2007) originally identified the H i cloud complex in a preliminary ALFALFA data set. These data have a channel width of 5 km s-1, a pixel size of 1′, and a spatial resolution of approximately 3.8′. We created our own version of the moment zero map of the H i emission in the cloud complex using the SoFiA (Serra et al., 2014, 2015) masking tool, adopting the standard smooth and clip algorithm with spatial smoothing kernels of 4.0 and 8.0′, with both no spectral smoothing and smoothing over two and four channels (10 and 20 km s-1). We set a detection threshold of 3.5 in order to include relatively faint emission, but coupled this with a high reliability threshold (95%) in order to ensure that only sources that are high S/N overall were retained. The moment zero map of the emission included in the SoFiA mask is shown in Figure 3 (left) as contours on a DECaLS optical image.
We estimate the H i mass of the entire structure by integrating all of the flux within the SoFiA mask and assuming a distance of 16.5 Mpc. This gives a value of . We note that this is about 0.1 dex higher than the value in Kent et al. (2009). However, in that work they considered each cloud component as a separate source, which may have led to some extended flux being missed. Whereas SoFiA’s spatial and spectral smoothing improves our ability to recover extended emission. In addition, Minchin et al. (2019) used deeper Arecibo observations to estimate a total H i mass of , arguing that the additional flux (compared to ALFALFA) comes from low column density, extended emission.666Minchin et al. (2019) also applied a 33% (0.1 dex) correction factor to account for the Arecibo/ALFA sidelobe response. Both Kent et al. (2009) and Minchin et al. (2019) adopt a distance of 16.7 Mpc for Virgo, which would increase their mass estimates by 2.4% (0.01 dex) relative to ours.
@@ 3.4 Hubble Space Telescope imaging
HST Advanced Camera for Surveys (ACS) images of BC6777http://dx.doi.org/10.17909/h12e-gb42 (catalog DOI: 10.17909/h12e-gb42) were taken in February 2023 as part of the program GO-17267 (PI: M. Jones). The imaging used the Wide Field Channel of ACS with the F606W and F814W filters. Over the course of two orbits, four exposures were taken in each filter, resulting in total integration times of approximately 2000 s in F606W and F814W. A false color composite of these images is shown in Figure 3 (right), centered on BC6.
@@ 3.5 Hobby-Eberly Telescope optical spectroscopy
We observed the main component of BC6 (the largest ellipse in Figure 3, right) with the Blue spectrograph of the Low Resolution Spectrograph 2 (LRS2-B, Chonis et al., 2016) on the 10m HET (Ramsey et al., 1998; Hill et al., 2021) under clear dark skies on the night of April 19th, 2023 with a 3300 s exposure. The LRS2-B spectrograph has and an integral field unit which is 6″12″ with 0.6″ lenslet fibers on sky, giving an abundance of sky fibers around the relatively compact optical emission from BC6. We use the H line to determine which spaxels include emission from BC6 and extract all spaxels within 1.5″ of the peak of the source with the LRS2Multi extension to the Panacea data reduction pipeline.888https://github.com/grzeimann/LRS2Multi and https://github.com/grzeimann/Panacea Briefly, this data reduction includes the necessary detector characterizations, fiber extraction, wavelength calibration, mirror illumination correction, sky subtraction, and 1-D spectral source extraction including a treatment of differential atmospheric refraction in the sky and object spectra. The resulting spectrum is shown in Figure 5.
@@
The spectrum is fairly typical of an H ii region, with clear, albeit relatively weak, emission lines. Only the H line is high enough S/N to be fit with a Gaussian. Thus, this line alone was used to determine the redshift km s-1 (with a barycentric correction applied). To obtain fluxes from the other lines we extract the integrated signal within a 2 region centered on the redshifted wavelength on each line. Here is the standard deviation of the Gaussian profile fit to the H line. Flux errors are estimated by summing the uncertainty values in the same wavelength range and dividing by the square root of the number of channels included.
@@
A reddening correction was calculated by assuming an intrinsic value of (for an H ii region at K; Osterbrock, 1989), an value of 3.1, and the Fitzpatrick (2004) extinction law. Based on the observed line ratio we calculated an extinction value of . This suggests that there is some internal extinction within BC6 as the Schlegel et al. (1998) dust map in this direction indicates minimal Galactic extinction (). Observed line fluxes were corrected for extinction (using the extinction model of Fitzpatrick, 2004) implemented in the Python package dust_extinction.999https://dust-extinction.readthedocs.io These fluxes and uncertainties are provided in Table 2. For consistency the quoted H flux does not correspond to that calculated with the Gaussian fit, but was instead calculated in the same manner as the other lines.
We note here that the extinction estimate is strictly speaking only applicable to the H ii region as it is likely the result of dust in its immediate vicinity. For simplicity we apply the Galactic extinction correction only for all stars in the HST CMD (§3.4 & Figure 4), even though for a few stars in the immediate vicinity of the H ii region the higher estimate is likely more appropriate.
@@
With the line flux measurements we used two line ratio metrics to estimate the gas-phase oxygen abundance in BC6 (following Pettini & Pagel, 2004), N2=[Nii]/H and O3N2=([Oiii]/H)/([Nii]/H). We note that these ratios are designed to be almost independent of our extinction correction as [Oiii]5007 and H occur at almost the same wavelength, as do [Nii]6583 and H. We see (in Table 2) that both metrics give almost exactly the same oxygen abundance, despite each method having an intrinsic scatter of 0.2 dex.
@@ 4.2 Metallicity
Following the same rationale as Jones et al. (2022b), the above metallicity estimate implies that the progenitor galaxy of BC6 was likely in the stellar mass range , assuming it roughly follows the mass–metallicity relation (Andrews & Martini, 2013). The broad range is a reflection of the large uncertainty in the metallicity. Unfortunately, over this stellar mass range the typical gas fraction () of a field galaxy smoothly transitions over the range (e.g. Huang et al., 2012), thus the fact that BC6 is accompanied by 109 M⊙ of H i gas does little to help narrow down the progenitor’s stellar mass. However, it does mean that the progenitor must have lost a significant fraction, probably the majority, of its H i gas reservoir.
@@ 4.3 Stellar mass
To estimate the stellar mass of BC6 we follow the same approach as Jones et al. (2022b), that is, assume a constant SFR and continuously build up a synthetic stellar population (of matching metallicity) until the observed total F814W magnitude of BC6 is reached. Using this approach we estimate a stellar mass of M⊙, which at the observed (NUV-based) SFR could have been built up over approximately 80 Myr. If we instead use and -band aperture photometry from the NGVS images and the scaling relations of Zibetti et al. (2009) and Taylor et al. (2011), we obtain estimates of and M⊙, respectively.
@@
With a stellar mass of M⊙, BC6 must have a gas fraction () of around 3,000 if we only consider the gas in the coincident cloud (C3; Figure 3, left), or over 20,000 if we include the whole H i cloud complex. This makes it one of, if not the most H i-rich extragalactic stellar systems known.
@@ 5.1 Point of origin
As mentioned by Minchin et al. (2019), tidal stripping does not seem to be a viable mechanism to form this structure. Not only is its isolation more difficult to explain with the low velocities (300 km s-1) involved in tidal stripping (e.g. Bournaud & Duc, 2006), but M⊙ represents a significant amount of H i gas for almost any galaxy and it is not possible to tidally remove a large fraction of a galaxy’s gas disk without also stripping stars. The exception to this would be if the progenitor was a particularly massive and H i-rich galaxy, however, that would also require a close passage from a similarly massive perturber to strip the gas, with neither remaining connected to the stripped gas in the present. Thus, the complete lack of stars over almost all of such a massive H i complex is a strong argument for ram pressure stripping and we will only consider the formation of BC6 via this mechanism.
@@
The H i cloud complex (Figure 3) is about 40′ across (190 kpc), making it plausible that the progenitor galaxy is a degree or more away. Unfortunately, despite there being no clear progenitor galaxy in the immediate vicinity of BC6, there are 10s of potential parent objects that are Virgo members within a few degrees (Figure 6). Both Kent et al. (2009) and Minchin et al. (2019) focused on NGC 4445 and NGC 4424 as these are only 30′ to the SW of BC6 and have similar radial velocities to the gas complex (346 km s-1 and 437 km s-1, respectively; Albareti et al., 2017; Haynes et al., 2018). NGC 4424 is in the process of being ram pressure stripped (Chung et al., 2009; Sorgho et al., 2017), but its H i tail points to the SE, roughly parallel to the H i complex containing BC6, rather than towards it. This makes NGC 4424 an unlikely point of origin of the H i complex. In the case of NGC 4445 the –B-band diameter scaling relation of Jones et al. (2018) indicates that its expected original H i content (before falling into the Virgo cluster) would have been (and it still contains ; Haynes et al., 2018). The observed mass of the H i complex therefore represents almost its entire original H i content. In Sorgho et al. (2017), figure A1, NGC 4445 also appears to have a ram pressure tail pointing away from the H i complex based on observations from the Karoo Array Telescope (KAT-7). However, Minchin et al. (2021) did not detect this tail in deep Arecibo observations, suggesting that it was likely an artifact in the KAT-7 data. Thus, NGC 4445 remains a possible point of origin, with the caveat that the H i mass of the complex represents virtually its entire original H i content.
@@
In simulations of high speed ram pressure stripping events (e.g. Kapferer et al., 2009; Tonnesen & Bryan, 2012, 2021; Göller et al., 2023), on scales much larger than a galaxy, the tails that are formed are usually relatively linear features. This means that for progenitor candidates that are within 1∘, but not aligned with the major axis of the H i complex, a larger than expected opening angle would be required to explain the geometry of the structure. It would perhaps still be possible to reproduce the observed geometry by involving clumpy structure within the ICM (the simulations assume the ICM is uniform and moving at a constant velocity). However, in an attempt to narrow our focus to the most likely candidates, we will now consider potential progenitors that are roughly NW and SE of BC6, such that the cloud geometry can readily be explained by a simple ram pressure stripping scenario. There are six potentially viable progenitors within a few degrees: NGCs 4380, 4390, and 4442 to the NW, and NGCs 4519, 4522, and 4535 to the SE. All are massive enough to meet the mass–metallicity requirement from §4.2, but three (NGCs 4380, 4442, and 4535) are actually above the expected mass range (stellar masses from Mosenkov et al., 2019). All, except NGC 4442, currently contain some H i, detected in ALFALFA. The basic properties of these galaxies are summarized in Table 3.
We start by considering NGC 4442 as this is the closest major galaxy in projection (and velocity) along the direction of the major axis of the H i complex. This galaxy is undetected in ALFALFA and its expected original H i content indicates that it has likely lost enough gas to account for the H i complex. However, this is also a relatively massive (; Mosenkov et al., 2019) and lenticular galaxy, and has probably been without H i gas for some time. It seems quite improbable that a major stripping event could have occurred and NGC 4442 could have changed from a gas-rich galaxy to a quenched lenticular with no signs of the interaction, in the time it took for the gas to move a few hundred kpc away. We therefore do not consider NGC 4442 a likely point of origin.
@@
We see from the metallicity measurements (Table 3) that most of the galaxies with measurements could plausibly be matches for BC6 (Table 2), with only NGC 4390 being significantly disfavored because of its metallicity. The predicted original H i mass of NGC 4390 also suggests that it never would have had enough H i to form the cloud complex associated with BC6. Furthermore, it appears NGC 4390 has not lost a significant amount of H i.101010We note that the predicted original H i mass is usually intended as a statistical quantity and has considerable uncertainty for any individual galaxy. For example, the measured H i content of NGC 4519 exceeds its predicted original H i mass by over 0.1 dex. These values should be considered a rough guide only.
@@
NGC 4380 was mapped in the VLA Imaging of Virgo in Atomic gas project (VIVA; Chung et al., 2009), which found that it has a truncated H i disk. This indicates a past episode of ram pressure stripping, however, the direction of stripping is unknown. This galaxy is also above the stellar mass range expected for the progenitor (; Mosenkov et al., 2019).
@@
Unfortunately, VIVA did not map either NGC 4519 or NGC 4535. Both are detected in ALFALFA, but these data have an angular resolution of 4′. At this resolution NGC 4535 appears undisturbed, but the H i distribution of NGC 4519 is clearly extended to the NW (towards BC6). However, NGC 4519 has a neighbor immediately to the NW (NGC 4519A) separated by only 200 km s-1 (Binggeli et al., 1985). Thus the disturbance in the H i distribution of NGC 4519 may be the result of either tides or ram pressure and with the resolution of the ALFALFA data there is little chance of distinguishing the two.